OpenStudy (anonymous):

determine the quadrants in which (x, y) is located so that x/y > 0

OpenStudy (anonymous):

\(\dfrac{x}{y}\) is positive if either both \(x\) and \(y\) are negative or positive. For example, \(x=y=-1\), you have \(\dfrac{-1}{-1}=1\). If \(x=y=1\), you have \(\dfrac{1}{1}=1\).

OpenStudy (anonymous):

in quad one: x and y are both positive in quad two: x is negative but y is positive in quad three: x is negative and y is negative in quad four: x is positive but y is negative so there